OpenStudy (anonymous):

solve the equation for y. 4x+7y-5x=8x-4+4y

OpenStudy (anonymous):

Arrange all y terms to the left side, the rest to the right side: Can you show it?

OpenStudy (anonymous):

how would i do that?

OpenStudy (anonymous):

Since 4y currently on the right side, move it to the left first

OpenStudy (anonymous):

so would i minus 4y to the 7y?

OpenStudy (anonymous):

then is comes out to 4x+3y-5x=8x-4 right?

OpenStudy (anonymous):

-1x?

OpenStudy (anonymous):

then move -x to the right side

OpenStudy (anonymous):

so i would add -x to 8x?

OpenStudy (anonymous):

when you switch the SIDE, then the SIGN changed! 8x + x = ...

OpenStudy (anonymous):

9x

OpenStudy (anonymous):

3y=9x-4?

OpenStudy (anonymous):

Final step is divide both side by 3 => y = ...

OpenStudy (anonymous):

@sonja_lee2 Where are you?

OpenStudy (anonymous):

y=9x-4/3?

OpenStudy (anonymous):

y=3x-1.333

OpenStudy (anonymous):

Actually, keep the fraction y = 3x - 4/3

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

can u also help me with solving for w. -3u-w=u+5w

OpenStudy (anonymous):

Just open the new post for each new question, so more helpers can solve, too :)

OpenStudy (anonymous):

k